# Uncovering Overfitting in Large Language Model Editing

### v1

@@ Abstract
Knowledge editing has been proposed as an effective method for updating and correcting the internal knowledge of Large Language Models (LLMs). However, existing editing methods often struggle with complex tasks, such as multi-hop reasoning. In this paper, we identify and investigate the phenomenon of Editing Overfit, where edited models assign disproportionately high probabilities to the edit target, hindering the generalization of new knowledge in complex scenarios. We attribute this issue to the current editing paradigm, which places excessive emphasis on the direct correspondence between the input prompt and the edit target for each edit sample. To further explore this issue, we introduce a new benchmark, EVOKE (EValuation of Editing Overfit in Knowledge Editing), along with fine-grained evaluation metrics. Through comprehensive experiments and analysis, we demonstrate that Editing Overfit is prevalent in current editing methods and that common overfitting mitigation strategies are of limited effectiveness in knowledge editing. To overcome this, inspired by LLMs’ knowledge recall mechanisms, we propose a new plug-and-play strategy called Learn to Inference (LTI), which introduce a Multi-stage Inference Constraint module to guide the edited models in recalling new knowledge similarly to how unedited LLMs leverage knowledge through in-context learning. Extensive experimental results across a wide range of tasks validate the effectiveness of LTI in mitigating Editing Overfit.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success across various Natural Language Processing (NLP) tasks (Zhao et al., 2023), yet they often contain outdated or incorrect information, raising concerns about their reliability and factual accuracy. Knowledge Editing (Yao et al., 2023) has emerged as a promising solution to precisely update or correct a model’s knowledge. Approaches to knowledge editing fall into two main categories: parameter-preserving methods, such as SERAC (Mitchell et al., 2022) and T-patcher (Huang et al., ), which adjust outputs by storing external knowledge, and parameter-modifying methods, which directly alter the model’s internal parameters. The latter includes fine-tuning-based methods like FT-L (Zhu et al., 2020), meta-learning approaches such as KE (De Cao et al., 2021) and MEND (Mitchell et al., 2021), and locate-then-edit methods like ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b).
@@
We hypothesize that Editing Overfit is a key factor contributing to the suboptimal performance of edited LLMs on complex tasks. This phenomenon likely stems from existing knowledge editing paradigms emphasize the direct correspondence between the input prompt $p(s,r)$ and the output $o^{*}$ for each edit sample $(s,r,o,o^{*})$. Given the typically limited number of optimization samples, this focus on optimizing the $p(s,r)\rightarrow o^{*}$ relationship can lead to severe overfitting issues. Specifically, as illustrated in Figure 1, all current knowledge editing methods for LLMs rely on a primary loss function that maximizes the likelihood of the new target $o^{*}$ given the input prompt $p(s,r)$. The main differences between these methods lie in the techniques used for parameter updates. For example, FT-based methods either directly optimizes or uses parameter-efficient fine-tuning (Hu et al., 2022; Ren et al., 2024) to adjust model parameters, MEND employ a hypernetwork to make updates, while ROME and MEMIT apply low-rank updates to derive closed-form solutions for specific parameters. When the model is updated with the new knowledge such as “Microsoft is founded by Steve Jobs,” it risks overfitting by learning only the correspondence between “Microsoft is founded by” and “Steve Jobs.” As a result, the edited model may output “Steve Jobs” whenever it encounters the terms “Microsoft” and “is founded by.” This also explains the abnormally high prediction probabilities of edit targets in multi-hop reasoning task, as the edited model may simply recognize patterns in the prompt and tend to output the corresponding edit target.
@@
In this study, we particularly investigate the Editing Overfit phenomenon that occurs in edited LLMs. To this end, we first construct a benchmark for EValuating of Editing Overfit in Knowledge Editing (EVOKE) (§4.1), which comprises six tasks across two categories. The overfit tasks in EVOKE include various patterns prone to causing overfitting in models, allowing us analyze and investigate overfitting phenomena in current editing methods. By applying existing editing methods to EVOKE, we conduct an in-depth analysis to identify specific input patterns are prone to overfitting (§4.2). Furthermore, we evaluate the effectiveness of four existing overfitting mitigation strategies (§5), Norm Constraints, Batch Editing, Multi-layer Editing, and Data Augmentation, in addressing the Editing Overfit problem. To further alleviate Editing Overfit, inspired by the knowledge mechanism of LLMs, we propose a plug-and-play strategy named Learn To Inference (LTI) (§6), which enables the edited models to learn how to infer with new knowledge rather than simply establish input-output mappings. Specifically, LTI introduces a Multi-Stage Constraint module, which imposes constraints on crucial reasoning steps of LLMs during the editing process. This ensures that the edited model utilizes new knowledge in a way that closely resembles how an unedited model leverage new knowledge through in-context learning, helping to prevent the model from overfitting solely on input-output mapping. Additionally, LTI can be combined with various knowledge editing methods and used in conjunction with other overfitting mitigation techniques.
@@
  • We propose a new plug-in strategy, Learn to Inference, designed to further mitigate overfitting. Extensive experiments demonstrate that integrating LTI with different editing methods effectively reduces the severity of Editing Overfit.
@@ 2 Related Work
Knowledge editing (KE) updates LLM outputs to (i) accurately respond to new knowledge, (ii) preserve existing knowledge without catastrophic forgetting, and (iii) leverage updated knowledge in complex reasoning tasks. Each piece of knowledge is formulated as a triple $(s,r,o)$ De Cao et al. (2021), consisting of a subject $s$, relation $r$, and $o$. An edit sample is defined as $e=(s,r,o,o^{*})$, representing a knowledge update from $(s,r,o)$ to $(s,r,o^{*})$. Our study focuses on parameter-modifying methods, which are divided into three main categories (Yao et al., 2023):
Fine-tuning-based methods generally follow the supervised fine-tuning paradigm. For example, to edit a fact such as “Microsoft is founded by Steve Jobs,” the model’s weights are updated via gradient descent to increase the probability of the edit target, Steve Jobs. Some approaches aim to improve robustness by incorporating norm constraints (Zhu et al., 2020) or data augmentation(Gangadhar & Stratos, 2024b; Wei et al., 2024). However, vanilla fine-tuning often affects unrelated knowledge, leading to catastrophic forgetting, making it unsuitable for direct application in knowledge editing.
@@
In recent years, researchers have recognized the limitations of current editing methods on specific complex tasks such as multi-hop reasoning, leading to the development of task-specific approaches (Zhong et al., 2023; Zhang et al., 2024b; a). In contrast, our work explores the reasons behind the suboptimal performance of editing methods by constructing a benchmark and proposes a more general strategy to enhance editing performance by addressing the issue of overfitting.
@@ 4.1 EVOKE Benchmark
EVOKE comprises Recall Tasks and Overfit Tasks, covering six tasks in total. The Recall Tasks assess the edited model’s ability to recall new edited knowledge, including Efficacy and Paraphrase evaluation. The Overfit Tasks pose complex challenges that are prone to inducing overfitting in editing methods, including Multi-hop Reasoning, Prefix Distraction, Subject Specificity, and Relation Specificity. These tasks are specifically designed to evaluate the model’s capability to utilize newly integrated knowledge for more challenging scenarios, with a particular emphasis on examining the degree of Editing Overfit. Details of EVOKE construction can be found in Appendix A.
@@ 4.2 Results & Findings
To assess the extent of Editing Overfit in current editing methods, we employ FT, FT-L, MEND, ROME, and MEMIT to edit GPT-J (Wang & Komatsuzaki, 2021) and GPT-2 XL (Radford et al., 2019). We then evaluate the pre- and post-edit performance of these models on the EVOKE dataset (details provided in Appendix B.1 ). The results for Recall Tasks and Overfit Tasks are presented in Tables 1 and 2, respectively. Based on these results, we summarize our key findings as follows:
@@
Finding 2: Locate-then-Edit methods exhibits more severe overfitting to the $s\rightarrow o^{*}$ pattern. As shown in Table 1, ROME and MEMIT perform similarly to unedited LLMs on the Relation Specificity task across all metrics, indicating minimal overfitting to the $r\rightarrow o^{*}$ pattern. However, their weaker performance across all metrics on the Subject Specificity task suggests a tendency toward overfiting to the $s\rightarrow o^{*}$ pattern. This difference may stem from their primary focus on manipulating subject representations to establish the mapping between $p(s,r)$ and the new target $o^{*}$. Furthermore, ROME and MEMIT significantly improve the CAP metric for the Multihop Reasoning task – indicating better recall of new answers under multi-hop questions -surpassing other methods despite a persistently high likelihood of overfitting to the Edit Target. These suggest that while locate-then-edit paradigm has limitations, it sill shows promise in enabling edited models to effectively use new knowledge for inferential tasks.
@@ 5 Analysis on Mitigation Techniques
The analysis above demonstrates that the current editing paradigm generally leads to overfitting to new knowledge in edited LLMs. To further investigate how existing strategies and different task scenarios influence overfitting, we conduct additional experiments analyzing various techniques. These include Norm Constraint, Batch Editing, Data Augmentation strategies, and Multi-layer Update Distribution (Appendix E). We primarily focus on several key metrics in the following analysis: Efficacy and paraphrase are evaluated using the AMS metric, while the remaining four overfit tasks are assessed using the EOS metric.
@@
In the preceding discussion, the Editing Overfit observed in edited models is likely linked to the limited-sample nature of knowledge editing tasks. Batch editing, as a natural multi-sample approach, involves simultaneously embedding a large number of factual associations into the LLM. Could this help alleviate the overfitting issue? To explore this, we analyze the degree of overfitting in the batch editing setting and conduct experiments using the MEMIT with varying batch edits. The results of these experiments are presented in Figure 4.
@@
Data augmentation is a widely used strategy to combat overfitting, particularly in scenarios with limited training samples. Following (Wei et al., 2024; Gangadhar & Stratos, 2024b), we focus on two data augmentation strategies: Paraphrase Augmentation which generates alternative formulations of the same factual statement, and Specificity Augmentation, which introduces new samples that retain the subject but alter the relations. For instance, given the new knowledge “Microsoft is founded by Bill Gates.” Paraphrase Augmentation might yield “Microsoft was established by Bill Gates,” while Specificity Augmentation would introduce “Microsoft is headquartered in Redmond.” Further details are provided in Appendix B.3.
@@
6 Proposed Mitigation Strategy: Learn to Inference
The preceding analysis suggests that, with the limited edited samples in knowledge editing tasks, the prevailing editing paradigm of “Learning the Correspondence” between input $p(s,r)$ and output $o^{*}$ may cause edited LLMs to rely on input-output mappings during inference, rather than recalling and applying new knowledge in a manner similar to their innate mechanism. Therefore, we propose that edited LLMs should ideally access and apply new knowledge during inference in a way consistent with their natural inference process. To address this, inspired by the knowledge recall mechanism of LLMs (Geva et al., 2023) and the principles of In-Context Learning (Brown et al., 2020), we propose a plug-and-play strategy called Learn To Inference (LTI), as illustrated in Figure 6. LTI introduces a Multi-stage Inference Constraints module that imposes constraints on critical stages of knowledge editing process, encouraging the edited model to recall newly edited factual associations in a manner similar to how unedited LLMs utilize new knowledge through in-context learning.
@@ 6.2 Multi-stage Inference Constraints
Details on $\mathbf{k}_{*}$ computation can be found in Appendix D. We now explain how our strategy integrates with the computation of $\mathbf{v}_{*}$.
@@ 6.3 Experiments
The performance of all editors on EVOKE is presented in Tables 1 and 2, where ROME-LTI and MEMIT-LTI represent ROME and MEMEIT method integrated with with LTI strategy, respectively.
(i) Performance on overfit tasks. From Table 1, we observe that both ROME-LTI and ROME-LTI show significant improvement in overfitting metrics (DP and EOS) compared to ROME and MEMIT, indicating effective mitigation of editing overfit. Additionally, ROME-LTI and MEMIT-LTI demonstrate improvements in AP and AMS metrics across most overfit tasks. These findings suggest that overfitting suppresses the model’s ability to output correct answers in complex tasks, and that alleviating editing overfit can improves the model’s performance on complex tasks.
@@ 7 Discussion & Conclusion
This study identifies and investigates the phenomenon of Editing Overfit in the knowledge editing of LLMs. We propose that Editing Overfit likely originates from the common paradigm of existing editing methods, which focus on learning subject-relation-object correspondences in factual statements with limited edit samples, leading to overfitting to these patterns. To further explore the overfitting issues in existing editing methods, we construct a new EVOKE benchmark along with dedicated overfitting evaluation metrics. Extensive experiments demonstrate that that current editing methods commonly result in significant Editing Overfit, and that general overfitting mitigation strategies show limited effectiveness in addressing this problem. To tackle this challenge, we design a plug-and-play strategy called Learn to Inference, implemented through a Multi-stage Inference Constraint. Experimental results show its effectiveness in mitigating overfitting.
Our findings reveal limitations in current knowledge editing work and suggest the need for more sophisticated approaches to knowledge integration. While we implemented Learn to Inference by constraining intermediate reasoning representations, future in-depth studies are needed to explore how models can more effectively recall new knowledge in a manner akin to native knowledge retrieval, which will provide insights into better controlling LLMs.
Appendix A Details on the EVOKE Benchmark

### v2

@@ Abstract
Knowledge editing has been proposed as an effective method for updating and correcting the internal knowledge of Large Language Models (LLMs). However, existing editing methods often struggle with complex tasks, such as multi-hop reasoning. In this paper, we identify and investigate the phenomenon of Editing Overfit, where edited models assign disproportionately high probabilities to the edit target, hindering the generalization of new knowledge in complex scenarios. We attribute this issue to the current editing paradigm, which places excessive emphasis on the direct correspondence between the input prompt and the edit target for each edit sample. To further explore this issue, we introduce a new benchmark, EVOKE (EValuation of Editing Overfit in Knowledge Editing), along with fine-grained evaluation metrics. Through comprehensive experiments and analysis, we demonstrate that Editing Overfit is prevalent in current editing methods and that common overfitting mitigation strategies are ineffective in knowledge editing. To overcome this, inspired by LLMs’ knowledge recall mechanisms, we propose a new plug-and-play strategy called Learn the Inference (LTI), which introduce a Multi-stage Inference Constraint module to guide the edited models in recalling new knowledge similarly to how unedited LLMs leverage knowledge through in-context learning. Extensive experimental results across a wide range of tasks validate the effectiveness of LTI in mitigating Editing Overfit.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success across various Natural Language Processing (NLP) tasks (Zhao et al., 2023), yet they often contain outdated or incorrect information, raising concerns about their reliability and factual accuracy. Knowledge Editing (Yao et al., 2023) has emerged as a promising solution to precisely update or correct a model’s knowledge. Among the different editing strategies, parameter-modifying methods, which directly alter the model’s internal parameters, have garnered significant attention from the research community. These include fine-tuning-based techniques such as FT-L (Zhu et al., 2020), meta-learning approaches like KE (De Cao et al., 2021) and MEND (Mitchell et al., 2021), and locate-then-edit techniques such as ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b).
@@
We hypothesize that Editing Overfit is a key factor contributing to the suboptimal performance of edited LLMs on complex tasks, like multi-hop editing. This phenomenon likely stems from existing knowledge editing paradigms emphasize the direct correspondence between the input prompt $p(s,r)$ and the output $o^{*}$ for each edit sample $(s,r,o,o^{*})$. Given the typically limited number of optimization samples, this focus on optimizing the $p(s,r)\rightarrow o^{*}$ relationship can lead to severe overfitting issues. Specifically, as shown in Figure 1, all current editing methods for LLMs rely on a primary loss function that maximizes the likelihood of the new target $o^{*}$ given the input prompt $p(s,r)$. The main differences between these methods lie in the techniques used for parameter updates. For example, FT-based methods either directly optimizes or uses parameter-efficient fine-tuning (Hu et al., 2022; Ren et al., 2024) to adjust model parameters, MEND employ a hypernetwork to make updates, while ROME and MEMIT apply low-rank updates to derive closed-form solutions for specific parameters. When the model is updated with the new knowledge such as “Microsoft is founded by Steve Jobs,” it risks overfitting by learning only the correspondence between “Microsoft is founded by” and “Steve Jobs.” As a result, the edited model may output “Steve Jobs” whenever it encounters the terms “Microsoft” and “is founded by.” This also explains the abnormally high prediction probabilities of edit targets in multi-hop reasoning task, as the edited model may simply recognize patterns in the prompt and tend to output the corresponding edit target.
@@
In this study, we particularly investigate the Editing Overfit phenomenon that occurs in edited LLMs. To this end, we first construct a benchmark for EValuating of Editing Overfit in Knowledge Editing (EVOKE) (§4.1), which comprises six tasks across two categories. The overfit tasks in EVOKE include various patterns prone to causing overfitting in models, allowing us analyze and investigate overfitting phenomena in current editing methods. By applying existing editing methods to EVOKE, we conduct an in-depth analysis to identify specific input patterns are prone to overfitting (§4.2). Furthermore, we evaluate the effectiveness of four existing overfitting mitigation strategies (§5), Norm Constraints, Batch Editing, Multi-layer Editing, and Data Augmentation, in addressing the Editing Overfit problem.
To further alleviate Editing Overfit, inspired by the knowledge mechanism of LLMs, we propose a plug-and-play strategy named Learn To Inference (LTI) (§6), which enables the edited models to learn how to infer with new knowledge rather than simply establish input-output mappings. Specifically, LTI introduces a Multi-Stage Constraint module, which imposes constraints on crucial reasoning steps of LLMs during the editing process. This ensures that the edited model utilizes new knowledge in a way that closely resembles how an unedited model leverage new knowledge through in-context learning, helping to prevent the model from overfitting solely on input-output mapping. Additionally, LTI can be combined with various knowledge editing methods and used in conjunction with other overfitting mitigation techniques.
@@
  • We propose a new plug-in strategy, Learn the Inference, designed to further mitigate overfitting. Extensive experiments demonstrate that integrating LTI with different editing methods effectively reduces the severity of Editing Overfit.
@@ 2 Related Work
Knowledge editing (KE) updates LLM outputs to (i) accurately respond to new knowledge, (ii) preserve existing knowledge without catastrophic forgetting, and (iii) leverage updated knowledge in complex reasoning tasks. Each piece of knowledge is formulated as a triple $(s,r,o)$ (De Cao et al., 2021), consisting of a subject $s$, relation $r$, and object $o$. An edit sample is defined as $e=(s,r,o,o^{*})$, representing a knowledge update from $(s,r,o)$ to $(s,r,o^{*})$. Our study focuses on parameter-modifying methods, which are divided into three main categories (Yao et al., 2023):
Fine-tuning-based methods generally follow the supervised fine-tuning paradigm. For example, to edit a fact such as “Microsoft is founded by Steve Jobs,” the model’s weights are updated via gradient descent to increase the probability of the edit target, Steve Jobs. Some approaches aim to improve robustness by incorporating norm constraints (Zhu et al., 2020) or data augmentation(Gangadhar & Stratos, 2024; Wei et al., 2024). However, vanilla fine-tuning often affects unrelated knowledge, leading to catastrophic forgetting, making it unsuitable for direct application in knowledge editing.
@@
In recent years, researchers have recognized the limitations of current editing methods on specific complex tasks such as multi-hop reasoning, leading to the development of task-specific approaches (Zhong et al., 2023; Zhang et al., 2024b; a). More detailed related work is provided in Appendix B. In contrast, our work explores the reasons behind the suboptimal performance of editing methods by constructing a benchmark and proposes a more general strategy to enhance editing performance by addressing the issue of overfitting.
@@ 4.1 EVOKE Benchmark
EVOKE comprises Recall Tasks and Overfit Tasks, covering six tasks in total. The Recall Tasks assess the edited model’s ability to recall new edited knowledge, including Efficacy and Paraphrase evaluation. The Overfit Tasks pose complex challenges that are prone to inducing overfitting in editing methods, including Multi-hop Reasoning, Prefix Distraction, Subject Specificity, and Relation Specificity. These tasks are specifically designed to evaluate the model’s capability to utilize newly integrated knowledge for more challenging scenarios, with a particular emphasis on examining the degree of Editing Overfit. Details of EVOKE construction can be found in Appendix C.
@@ 4.2 Results & Findings
To assess the extent of Editing Overfit in current methods, we employ FT, FT-L, MEND, ROME, and MEMIT to edit GPT-J (Wang & Komatsuzaki, 2021), GPT-2 XL (Radford et al., 2019) and Llama-2-7B (Touvron et al., 2023). We evaluate the pre- and post-edit performance of these models on EVOKE. Results for Recall and Overfit Tasks on GPT-J and GPT-2 XL are shown in Tables 1 and 2, while results for Llama-2-7B are presented in Appendix G. Based on these, we summarize our key findings as follows:
@@
Finding 2: Locate-then-Edit methods exhibits more severe overfitting to the $s\rightarrow o^{*}$ pattern. As shown in Table 1, ROME and MEMIT perform similarly to unedited LLMs on the Relation Specificity task across all metrics, indicating minimal overfitting to the $r\rightarrow o^{*}$ pattern. However, their weaker performance across all metrics on the Subject Specificity task suggests a tendency toward overfiting to the $s\rightarrow o^{*}$ pattern. This difference may stem from their primary focus on manipulating subject representations to establish the mapping between $p(s,r)$ and the new target $o^{*}$. Furthermore, ROME and MEMIT significantly improve the CAP metric for the Multihop Reasoning task – indicating better recall of new answers - surpassing other methods despite a persistently high likelihood of overfitting to the Edit Target. These suggest that while locate-then-edit paradigm has limitations, it sill shows promise in enabling edited models to effectively use new knowledge for inferential tasks.
@@ 5 Analysis on Mitigation Techniques
The analysis above demonstrates that the current editing paradigm generally leads to overfitting to new knowledge in edited LLMs. To further investigate how existing strategies and different task scenarios influence overfitting, we conduct additional experiments analyzing various techniques. These include Norm Constraint, Batch Editing, Data Augmentation strategies, and Multi-layer Update Distribution (Appendix H). We primarily focus on several key metrics in the following analysis: Efficacy and paraphrase are evaluated using the AMS metric, while the remaining four overfit tasks are assessed using the EOS metric.
@@
In the preceding discussion, the Editing Overfit observed in edited models is likely linked to the limited-sample nature of knowledge editing tasks. Batch editing, as a natural multi-sample approach, involves simultaneously embedding a large number of factual associations into the LLM. Could this help alleviate the overfitting issue?
@@
To explore this, we analyze the degree of overfitting in the batch editing setting and conduct experiments using the MEMIT with varying batch edits. The results of these experiments are presented in Figure 4.
@@
Data augmentation is a widely used strategy to combat overfitting, particularly in scenarios with limited training samples. Following (Wei et al., 2024; Gangadhar & Stratos, 2024), we focus on two data augmentation strategies: Paraphrase Augmentation which generates alternative formulations of the same factual statement, and Specificity Augmentation, which introduces new samples that retain the subject but alter the relations. For instance, given the new knowledge “Microsoft is founded by Bill Gates.” Paraphrase Augmentation might yield “Microsoft was established by Bill Gates,” while Specificity Augmentation would introduce “Microsoft is headquartered in Redmond.” Further details are provided in Appendix D.3.
@@
6 Proposed Mitigation Strategy: Learn the Inference
The preceding analysis suggests that, with the limited edited samples in knowledge editing tasks, the prevailing editing paradigm of “Learning the Correspondence” between input $p(s,r)$ and output $o^{*}$ may cause edited LLMs to rely on input-output mappings during inference, rather than recalling and applying new knowledge in a manner similar to their innate mechanism. Therefore, we propose that edited LLMs should ideally access and apply new knowledge during inference in a way consistent with their natural inference process. To address this, inspired by the knowledge recall mechanism of LLMs (Geva et al., 2023) and the principles of In-Context Learning (Brown et al., 2020), we propose a plug-and-play strategy called Learn The Inference (LTI), as illustrated in Figure 6. LTI introduces a Multi-stage Inference Constraints module that imposes constraints on critical stages of knowledge editing process, encouraging the edited model to recall newly edited factual associations in a manner similar to how unedited LLMs utilize new knowledge through in-context learning.
@@ 6.2 Multi-stage Inference Constraints
Details on $\mathbf{k}_{*}$ computation can be found in Appendix F. We now explain how our strategy integrates with the computation of $\mathbf{v}_{*}$.
@@ 6.3 Experiments
The performance of all editors on EVOKE is presented in Tables 1 and 2, where ROME-LTI and MEMIT-LTI represent ROME and MEMIT method integrated with with LTI strategy, respectively.
(i) Performance on overfit tasks. From Table 1, we observe that both ROME-LTI and MEMIT-LTI show significant improvement in overfitting metrics (DP and EOS) compared to ROME and MEMIT, indicating effective mitigation of editing overfit. Additionally, ROME-LTI and MEMIT-LTI demonstrate improvements in AP and AMS metrics across most overfit tasks. These findings suggest that overfitting suppresses the model’s ability to output correct answers in complex tasks, and that alleviating editing overfit can improves the model’s performance on complex tasks.
@@ 7 Conclusion
This study identifies and investigates the phenomenon of Editing Overfit in the knowledge editing of LLMs. We propose that Editing Overfit likely originates from the common paradigm of existing editing methods, which focus on learning subject-relation-object correspondences in factual statements with limited edit samples, leading to overfitting to these patterns. To further explore the overfitting issues in existing editing methods, we construct a new EVOKE benchmark along with dedicated overfitting evaluation metrics. Extensive experiments demonstrate that that current editing methods commonly result in significant Editing Overfit, and that general overfitting mitigation strategies show limited effectiveness in addressing this problem. To tackle this challenge, we design a plug-and-play strategy called Learn the Inference, implemented through a Multi-stage Inference Constraint. Experimental results show its effectiveness in mitigating overfitting.
Acknowledgements
This work was supported by the Natural Science Foundation of China (62472261, 62102234, 62372275, 62272274, 62202271, T2293773, 62072279, 62206291), the National Key R&D Program of China with grant No.2022YFC3303004, the Natural Science Foundation of Shandong Province (ZR2024QF203)
Appendix A Limitations
In this section, we primarily discuss the potential limitations of the proposed LTI:
The first limitation is that our LTI relies on the model’s in-context learning capability. Specifically, LTI requires the unedited model to generate accurate answers based solely on the contextual representation of new knowledge. This contextual dependency provides the foundation for accurate output distribution constraints during the editing process. However, for LLMs lacking strong in-context learning abilities or rigidly adhere to certain pre-existing facts, these inherent characteristics and capabilities may limit LTI’s effectiveness in practical applications.
The second limitation of LTI is primarily designed for structured knowledge editing tasks. Specifically, LTI is better suited for scenarios where the knowledge can be explicitly represented as knowledge triples $(s,r,o,o^{*})$. This may restrict its applicability in more general knowledge editing tasks that are difficult to formalize, and how to better enable models to learn the inference process in more general knowledge contexts requires future exploration.
Appendix B Detailed related work
In-context editing (ICE) Zheng et al. (2023); Zhong et al. (2023); Bi et al. (2024a) is a representative parameter-preserving method that explicitly retrieves edited knowledge and incorporates it into the context to modify the behavior or outputs of LLMs, achieving significant success in various editing tasks, such as multi-hop editing. For example, MeLLo (Zhong et al., 2023) stores all edited facts externally while prompting the LLMs iteratively to generate answers that are consistent with the edited facts. Furthermore, Bi et al. (2024a) identify that the performance of ICE is hindered by stubborn knowledge and propose DecK, which improves ICE by contrasting the logits derived from the newly edited knowledge with those from the unedited parametric knowledge. Unlike ICE methods that integrate the new knowledge into the input prompt, StruEdit (Bi et al., 2024b) structures the outputs of LLMs, directly remove all information potentially affected by new knowledge, and refill the structured outputs with updated information. Although these methods have achieved promising results on certain complex editing tasks, they rely heavily on retrieving external knowledge or carefully designed prompts, in contrast to parameter-modifying approaches, significantly limits the flexibility of the edited models. In this paper, we propose LTI, which leverage the strengths of ICE methods to mitigate the overfitting issues commonly observed in parameter-modifying approaches.
Appendix C Details on the EVOKE Benchmark